# magnetohydrodynamics (MHD) — turbulence — diffusion — protoplanetary disks

### v1

@@ 3.2.3 Disappearance of the Vertical Magnetic Flux in the Transition Zone
In this section, we investigate how the vertical flux disappears in the transition zone (Figures 7d and 9d). This flux transport is very efficient, and occurs in less time than 7 rotations at $R=R_{\mathrm{MRI}}$.
@@
The evolution of the vertical field is determined by Equation (40). In the early evolution, $\langle(E_{\phi})_{\mathrm{O}}\rangle$ provides a dominant contribution to $\langle E_{\phi}\rangle$ as will be shown in Figure 18. Because $|\partial B_{R}/\partial z|$ dominates over $|\partial B_{z}/\partial R|$
@@
Finally, we investigate the radial transport of the vertical field. Equation (40) shows that the transport velocity of $\langle B_{z}\rangle$ is estimated by $\langle E_{\phi}\rangle/\langle B_{z}\rangle$. Since $\langle B_{z}\rangle$ is positive in most regions at the mid-plane, the sign of $\langle E_{\phi}\rangle$ indicates the direction of the vertical field transport. Figure 18 shows the time evolution of the radial profiles of the toroidal electric fields at the mid-plane. OR mainly contributes to $\langle E_{\phi}\rangle$ while $\langle E_{\phi}\rangle_{\mathrm{I}}$ is negligible at the mid-plane where OR suppresses the MRI. AD plays a minor role in $\langle E_{\phi}\rangle$ except for $R\gtrsim 1~{}\mathrm{au}$. A sharp increase in $\langle E_{\phi}\rangle$ in $R\lesssim 0.9~{}\mathrm{au}$ is attributed to a strong radial dependence of $\langle\eta_{\mathrm{O}}\rangle$ in $R<R_{\eta}$ (Figure 2). The radial gradient of $R\langle E_{\phi}\rangle$ around $R\sim 0.8~{}\mathrm{au}$ increases with time from $t\sim 50~{}t_{\mathrm{K0}}$ to $70~{}t_{\mathrm{K0}}$, representing the rapid outward transport of $\langle B_{z}\rangle$ (Figure 16). The rapid outward transport of the vertical field in the transition zone is due to steepening of the vertical gradient of $B_{R}$ around the mid-plane by AD as seen in Figures 16 and 17.
@@
The turbulence and dynamo activity in the active zone drives time variation in the transition and coherent zones. Figure 19 shows the radius-time diagrams of $\langle B_{z}\rangle_{H}$ and $\langle B_{\phi}\rangle_{H}$. As shown in Section 3.2.1, in the active zone ($R\leq R_{\mathrm{MRI}}$), the ring and gap structures develop in $\langle B_{z}\rangle_{H}$ and $\langle\rho\rangle_{H}$.
@@ 4.1 Mass and Angular Momentum Transfer Throughout the Disk
In this section, we investigate the radial dependence of the mass accretion rate throughout the disk. The mass accretion rate at a given $R$ is defined as
@@
Figure 23: Radial dependence of the mass accretion rate in the early stage for HighRes (left) and LowRes (right). All the quantities shown in this figure are averaged over $400t_{\mathrm{K0}}\leq t\leq 500t_{\mathrm{K0}}$. The first row compares the mass accretion rate (pink) integrated over $|z|\leq H$ and their predictions (black) given by Equation (53). The blue and orange lines correspond to $\dot{M}_{R\phi}^{H}$ and $\dot{M}_{\phi z}^{H}$, respectively. The second row is the same as the first row but the mass accretion rate is integrated over $|z|\leq z_{\mathrm{atm}}$. The radial profiles of the surface densities are shown in the third row. In the fourth row, the color maps of $R^{2}\langle\rho v_{R}\rangle_{t}$ are displayed in the plane of ($\log_{10}R,z/H$), and the black solid and dashed lines correspond to $z=\pm H$ and $z=\pm z_{\mathrm{atm}}$. In each panel, the two vertical dotted lines show $R=R_{\mathrm{MRI}}$ and $R=R_{\eta}$. In the first and second row, to reduce significant spatial fluctuations of the mass accretion rates $\dot{M}_{R\phi}^{H,z_{\mathrm{atm}}}$, $\dot{M}_{\phi z}^{H,z_{\mathrm{atm}}},\dot{M}_{\mathrm{pred}}^{H,z_{\mathrm{atm}}}$, the centered moving averages over $\Delta R/R=0.4$ are taken. All the quantities are shown in the code units.
@@ 4.1.1 Early Evolution
The top and second rows of Figure 23 show comparison between the measured and predicted mass accretion rates by integrating over $|z|\leq H$ and $|z|\leq z_{\mathrm{atm}}$, respectively. Equation (53) reproduces the actual mass accretion rates reasonably well for all the panels.
Next, we investigate the mass transfer in the active zone around the mid-plane ($|z|\leq H$). Figures 23a and 23b show that the mass transfer is mainly driven by the $R\phi$ torque due to the MRI turbulence while the contribution from the $\phi z$ torque is negligible. The active zone shows that the radial mass fluxes are negative at $R\lesssim 0.25~{}$au while they are positive in the outer region. This is a typical feature of viscous evolution of an isolated accretion disk where the outer region expands outward by receiving the angular momentum from the inner region where the gas accretes inwards. The resolution dependence of the radial mass flux is seen especially in the expanding outer region ($0.25~{}\mathrm{au}\lesssim R\lesssim R_{\mathrm{MRI}}$). The outward radial mass flux around $R\sim 0.4~{}\mathrm{au}$ is more significant in HighRes than in LowRes run.
When the upper layers of the disk is considered in the estimation of the mass accretion rate, the resolution dependence becomes more significant. For HighRes run, the mass accretion rate is not influenced by the upper layers significantly, or $\dot{M}^{H}\sim\dot{M}^{z_{\mathrm{atm}}}$ (Figures 23a and 23c). By contrast, LowRes run shows that $\langle W_{\phi z}\rangle$ owing to large-scale magnetic fields in the upper layers makes $\dot{M}_{z_{\mathrm{atm}}}$ positive in the inner active zone although $\dot{M}^{H}$ is negative there. The corresponding structure in Figure 23h is the outgoing gas upper layers in the active zone. This is because the magnetic fields are more coherent for LowRes run than HighRes run.
@@ 4.1.2 Later Evolution
Figure 25: The same as Figure 23 but the quantities are averaged over $1500\leq t/t_{\mathrm{K0}}\leq 2000$ for LowRes run. In Panels (a) and (b), to reduce significant spatial fluctuations of the mass accretion rates $\dot{M}_{R\phi}^{H,z_{\mathrm{atm}}}$, $\dot{M}_{\phi z}^{H,z_{\mathrm{atm}}},\dot{M}_{\mathrm{pred}}^{H,z_{\mathrm{atm}}}$, the centered moving averages over $\Delta R/R=0.4$ are taken. All the quantities are shown in the code units.
Unlike in the early evolution, the outer regions of the active zone ($R\sim 0.4~{}\mathrm{au}$) do not show outward mass flux $\dot{M}^{H}>0$ and $\dot{M}^{z_{\mathrm{atm}}}>0$ in the long-term evolution (Figures 25a and 25b). The net gas accretion flows through the disk with $|z|\leq H$ and $|z|\leq z_{\mathrm{atm}}$ are almost zero. One can identify the gas flow converging to the rings both in Figures 25a and 25b. As will be shown in Section 4.2.1, gas accretion in higher latitudes to the inner boundary continues in the late evolution.
The behaviors around the inner edge of the transition zone shown in Figure 25 are similar to those shown in Figure 23. Inside the disk ($R\sim R_{\mathrm{MRI}}$, $|z|\leq H$), the mass is transferred inward by the $\phi z$ torque due to the hour-glass-shaped magnetic fields. When the upper layers of the disk ($|z|\leq z_{\mathrm{atm}}$) is considered, the inward angular momentum transfer is almost compensated by the outward angular momentum transfer owing to the $R\phi$ torque. In Section 4.2, we will see that the absence of net angular momentum transport around the inner edge of the transition zone results in almost zero radial mass transfer from the active zone to the transition zone (Figure 26).
@@
The mass accretion rate in the coherent zone is determined by $\langle W_{\phi z}\rangle$. Using the ratio $f_{\phi z}=B_{\phi}/B_{z}$ and Equation (55), the mass accretion rate driven by magnetic braking is estimated to be
$\dot{M}=\frac{8\pi R}{\Omega_{\mathrm{K}}}\frac{f_{\phi z}B_{z}^{2}}{4\pi}=2\pi R\Sigma c_{\mathrm{s,mid}}\times\left(4\sqrt{2/\pi}f_{\phi z}\beta_{z}^{-1}\right),$ (57)
where $\beta_{z}=8\pi\rho_{\mathrm{mid}}c_{\mathrm{s,mid}}^{2}/B_{z}^{2}$. Equation (57) predicts that $\dot{M}\sim 10^{-3}$ in the code units at $R=1~{}\mathrm{au}$ and $\dot{M}$ decreases in proportion to $\propto R^{-1/4}$ since $\Sigma\propto R^{-1}$ (Equation (5)), where we use the fact that the radial magnetic flux transport from the transition zone decreases $\beta_{z}$ in the coherent zone by a factor of two from the initial plasma beta $10^{4}$ (Section 3.2.3) and $f_{\phi z}$ is estimated to $\sim 5$ from the simulation result.
@@ 4.2 Time Evolution of the Masses of the Three Zones
Figure 26: Mass transfer rate through the four surfaces of (a) the active, (b) transition, and (c) coherent zones for LowRes run. The blue, orange, and green lines correspond to $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{R+}$, $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{R-}$, and $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{z}$, respectively. The gray lines show $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{\mathrm{tot}}$. To reduce significant temporal fluctuations of the mass transfer rates, we take the $20t_{\mathrm{K0}}$ centered moving average for them. All the quantities are shown in the code units.
@@
where the subscripts ”ac”, ”tr”, and ”co” stand for the active, transition, and coherent zones, respectively, the mass increasing rates passing through the surfaces $R=R_{-}$ and $R_{+}$ are denoted by $(\dot{M}_{\mathrm{ac,tr,co}})_{R-}$ and $(\dot{M}_{\mathrm{ac,tr,co}})_{R+}$, respectively, and the mass increasing rates from the surfaces $z=-z_{\mathrm{atm}}$ and $z=+z_{\mathrm{atm}}$ are added and denoted as $(\dot{M}_{\mathrm{ac,tr,co}})_{z}$. The signs of $(\dot{M}_{\mathrm{ac,tr,co}})_{R-}$, $(\dot{M}_{\mathrm{ac,tr,co}})_{R+}$, and $(\dot{M}_{\mathrm{ac,tr,co}})_{z}$ are determined so that they are positive when the total mass inside the enclosed region increases owing to its contribution. Since the surface at $R=R_{\mathrm{MRI}}$ ($R=R_{\eta}$) is shared between the active and transition zones (the transition and coherent zones), $(\dot{M}_{\mathrm{ac}})_{R+}=-(\dot{M}_{\mathrm{tr}})_{R-}$ ($(\dot{M}_{\mathrm{tr}})_{R+}=-(\dot{M}_{\mathrm{co}})_{R-}$) is satisfied. Figure 26 shows the time evolution of $(\dot{M}_{\mathrm{ac,tr,co}})_{R+}$, $(\dot{M}_{\mathrm{ac,tr,co}})_{R-}$ , $(\dot{M}_{\mathrm{ac,tr,co}})_{z}$, and $(\dot{M}_{\mathrm{ac,tr,co}})_{\mathrm{tot}}$.

### v2

@@ 3.2.3 Disappearance of the Vertical Magnetic Flux in the Transition Zone
In this section, we investigate how the vertical flux disappears in the transition zone (Figures 7d and 9d). We will present that both a steep spatial gradient of $\eta_{\mathrm{O}}$ and amplification of the toroidal electric field by AD are necessary for efficient outward transfer of the vertical flux. This flux transport is very efficient, and occurs in less time than 7 rotations at $R=R_{\mathrm{MRI}}$.
@@
We investigate the radial transport of the vertical field. Equation (40) shows that the radial transport velocity of $\langle B_{z}\rangle$ is estimated by $\langle E_{\phi}\rangle/\langle B_{z}\rangle$. Since $\langle B_{z}\rangle$ is positive in most regions at the mid-plane, the sign of $\langle E_{\phi}\rangle$ represents the direction of the vertical field transport. Figure 17 shows the time evolution of the radial profiles of the toroidal electric fields at the mid-plane. OR mainly contributes to $\langle E_{\phi}\rangle$ while $\langle E_{\phi}\rangle_{\mathrm{I}}$ is negligible at the mid-plane where OR suppresses the MRI. AD plays a minor role in $\langle E_{\phi}\rangle$ except for $R\gtrsim 1~{}\mathrm{au}$. A sharp increase in $\langle E_{\phi}\rangle$ in $R\lesssim 0.9~{}\mathrm{au}$ is attributed to the strong radial dependence of $\langle\eta_{\mathrm{O}}\rangle$ in $R<R_{\eta}$ (Figure 2). The radial gradient of $R\langle E_{\phi}\rangle$ around $R\sim 0.8~{}\mathrm{au}$ increases with time from $t\sim 50~{}t_{\mathrm{K0}}$ to $70~{}t_{\mathrm{K0}}$, representing the rapid outward transport of $\langle B_{z}\rangle$ (Figure 16c).
@@
Figure 18 shows that an increase in $E_{\phi}$ results in outward magnetic flux transfer. We investigate how $\langle E_{\phi}\rangle_{\mathrm{O}}$ is generated. Because $|\partial B_{R}/\partial z|$ dominates over $|\partial B_{z}/\partial R|$
@@
In this section, we will show that the MRI activity drives quasi-periodic variations of the magnetic fields in the transition and coherent zones. The sign of $B_{\phi}$ reverses quasi-periodically. Anti-diffusion caused by AD generates a $B_{z}$ condensation where the sign of $B_{\phi}$ is flipped. The magnetic field disturbances do not propagate beyond $R=R_{\mathrm{ch}}$.
Figure 19 shows the radius-time diagrams of $\langle B_{z}\rangle_{H}$ and $\langle B_{\phi}\rangle_{H}$. As shown in Section 3.2.1, in the active zone ($R\leq R_{\mathrm{MRI}}$), the ring and gap structures develop in $\langle B_{z}\rangle_{H}$ and $\langle\rho\rangle_{H}$.
@@ 4.1 Mass and Angular Momentum Transfer Throughout the Disk
In this section, we investigate the radial dependence of the mass accretion rate throughout the disk. The mass transfer rate at a given $R$ is defined as
@@
Figure 23: Radial dependence of the mass transfer rate in the early stage for HighRes (left) and LowRes (right). All the quantities shown in this figure are averaged over $400t_{\mathrm{K0}}\leq t\leq 500t_{\mathrm{K0}}$. The first row compares the mass transfer rate (pink) integrated over $|z|\leq H$ and their predictions (black) given by Equation (53). The blue and orange solid lines correspond to $\dot{M}_{R\phi}^{H}$ and $\dot{M}_{\phi z}^{H}$, respectively. The blue and orange dashed lines represent $\dot{M}_{R\phi}^{H}$ and $\dot{M}_{\phi z}^{H}$ evaluated without the Reynolds stress, respectively. The second row is the same as the first row but the mass transfer rate is integrated over $|z|\leq z_{\mathrm{atm}}$. The radial profiles of the surface densities are shown in the third row. In the fourth row, the color maps of $R^{2}\langle\rho v_{R}\rangle_{t}$ are displayed in the plane of ($\log_{10}R,z/H$), and the black solid and dashed lines correspond to $z=\pm H$ and $z=\pm z_{\mathrm{atm}}$. In each panel, the two vertical dotted lines show $R=R_{\mathrm{MRI}}$ and $R=R_{\eta}$. In the first and second row, to reduce significant spatial fluctuations of the mass transfer rates $\dot{M}_{R\phi}^{H,z_{\mathrm{atm}}}$, $\dot{M}_{\phi z}^{H,z_{\mathrm{atm}}},\dot{M}_{\mathrm{pred}}^{H,z_{\mathrm{atm}}}$, the centered moving averages over $\Delta R/R=0.4$ are taken. All the quantities are shown in the code units.
@@ 4.1.1 Early Evolution
Mass Transfer in the Active Zone
The top and second rows of Figure 23 show comparison between the measured and predicted mass transfer rates by integrating over $|z|\leq H$ and $|z|\leq z_{\mathrm{atm}}$, respectively. Equation (53) reproduces the actual mass transfer rates reasonably well for all the panels.
Next, we investigate the mass transfer in the active zone around the mid-plane ($|z|\leq H$). Figures 23a and 23b show that the mass transfer is mainly driven by the $R\phi$ torque due to the MRI turbulence while the contribution from the $\phi z$ torque is negligible. For both torques, the Maxwell stress plays a more dominant role than the Reynolds stress does. The active zone shows that the radial mass fluxes are negative at $R\lesssim 0.25~{}$au while they are positive in the outer region. This is a typical feature of viscous evolution of an isolated accretion disk where the outer region expands outward by receiving the angular momentum from the inner region where the gas accretes inwards. The resolution dependence of the radial mass flux is seen especially in the expanding outer region ($0.25~{}\mathrm{au}\lesssim R\lesssim R_{\mathrm{MRI}}$). The outward radial mass flux around $R\sim 0.4~{}\mathrm{au}$ is more significant in HighRes than in LowRes run.
When the upper layers of the disk is considered in the estimation of the mass transfer rate, the resolution dependence becomes more significant. For HighRes run, the mass transfer rate is not influenced by the upper layers significantly, or $\dot{M}^{H}\sim\dot{M}^{z_{\mathrm{atm}}}$ (Figures 23a and 23c). By contrast, LowRes run shows that $\langle W_{\phi z}\rangle$ owing to large-scale magnetic fields in the upper layers makes $\dot{M}_{z_{\mathrm{atm}}}$ positive in the inner active zone although $\dot{M}^{H}$ is negative there. The corresponding structure in Figure 23h is the outgoing gas upper layers in the active zone. This is because the magnetic fields are more coherent for LowRes run than HighRes run.
@@
Mass Transfer Around the Inner Edge of the Transition Zone
@@
Ring Structures Around the Inner Edge of the Transition Zone
@@ 4.1.2 Later Evolution
Figure 25: The same as Figure 23 but the quantities are averaged over $1500\leq t/t_{\mathrm{K0}}\leq 2000$ for LowRes run. In Panel (b), the mass transfer rate integrated over $|z|\leq 5z_{\mathrm{atm}}$ is plotted. In Panels (a) and (b), to reduce significant spatial fluctuations of the mass transfer rates $\dot{M}_{R\phi}^{H,z_{\mathrm{atm}}}$, $\dot{M}_{\phi z}^{H,z_{\mathrm{atm}}},\dot{M}_{\mathrm{pred}}^{H,z_{\mathrm{atm}}}$, and $\dot{M}^{H,z_{\mathrm{atm}},5z_{\mathrm{atm}}}$, the centered moving averages over $\Delta R/R=0.4$ are taken. All the quantities are shown in the code units.
Unlike in the early evolution, the outer regions of the active zone ($R\sim 0.4~{}\mathrm{au}$) do not show outward mass flux $\dot{M}^{H}>0$ and $\dot{M}^{z_{\mathrm{atm}}}>0$ in the long-term evolution (Figures 25a and 25b). The net mass transfer rates through the disk with $|z|\leq H$ and $|z|\leq z_{\mathrm{atm}}$ are almost zero. One can identify the gas flow converging to the rings both in Figures 25a and 25b.
The fact that gas accretion does not occur around the disk mid-plane may be due to the ring structures. Jacquemin-Ide et al. (2021) investigated an MRI disk (model SEp) similar to our active zone and found that the radial migration of the ring structures is not seen. However, the physical mechanism is unclear and needs to be investigated. We note that the insufficient resolution of LowRes run may cause the almost zero accretion rate. The gas accretion continues in the higher latitudes although it halts inside the disk. This is evident that the mass transfer rate integrated over $|z|\leq 5z_{\mathrm{atm}}$ ($\dot{M}^{5z_{\mathrm{atm}}}$) takes negative values except around the rings (also see Figure 27).
The behaviors around the inner edge of the transition zone shown in Figure 25 are similar to those shown in Figure 23. Inside the disk ($R\sim R_{\mathrm{MRI}}$, $|z|\leq H$), the mass is transferred inward by the $\phi z$ torque due to the hour-glass-shaped magnetic fields. When the upper layers of the disk ($|z|\leq z_{\mathrm{atm}}$) is considered, the inward angular momentum transfer is almost compensated by the outward angular momentum transfer owing to the $R\phi$ torque. In Section 4.2, we will see that the absence of net angular momentum transport around the inner edge of the transition zone results in almost zero radial mass transfer from the active zone to the transition zone (Figures 26a and 26c).
@@
The mass transfer rate in the coherent zone is determined by $\langle W_{\phi z}\rangle$. Using the ratio $f_{\phi z}=B_{\phi}/B_{z}$ and Equation (55), the mass transfer rate driven by magnetic braking is estimated to be
$\dot{M}=-\frac{8\pi R}{\Omega_{\mathrm{K}}}\frac{f_{\phi z}B_{z}^{2}}{4\pi}=-2\pi R\Sigma c_{\mathrm{s,mid}}\times\left(4\sqrt{2/\pi}f_{\phi z}\beta_{z}^{-1}\right),$ (57)
where $\beta_{z}=8\pi\rho_{\mathrm{mid}}c_{\mathrm{s,mid}}^{2}/B_{z}^{2}$. Equation (57) predicts that $\dot{M}\sim-10^{-3}$ in the code units at $R=1~{}\mathrm{au}$ and $|\dot{M}|$ decreases in proportion to $\propto R^{-1/4}$ since $\Sigma\propto R^{-1}$ (Equation (5)), where we use the fact that the radial magnetic flux transport from the transition zone decreases $\beta_{z}$ in the coherent zone by a factor of two from the initial plasma beta $10^{4}$ (Section 3.2.3) and $f_{\phi z}$ is estimated to $\sim 5$ from the simulation result.
@@ 4.2 Time Evolution of the Masses of the Three Zones
Figure 26: Mass transfer rate through the four surfaces of (a) the active, (c) transition, and (d) coherent zones for LowRes run. The blue, orange, and green lines correspond to $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{R+}$, $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{R-}$, and $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{z}$, respectively. The gray lines show $\left(\dot{M}_{\mathrm{ac,tr,co}}\right)_{\mathrm{tot}}$. To reduce significant temporal fluctuations of the mass transfer rates, we take the $20t_{\mathrm{K0}}$ centered moving average for them. Panel (b) shows the time variations of $R\langle B_{\phi}\rangle$ at $z=2H$ and two different radii ($R=0.3~{}$au and $R_{\mathrm{MRI}}=0.5$ au). All the quantities are shown in the code units.
@@
where the subscripts ”ac”, ”tr”, and ”co” stand for the active, transition, and coherent zones, respectively, the mass increasing rates passing through the surfaces $R=R_{-}$ and $R_{+}$ are denoted by $(\dot{M}_{\mathrm{ac,tr,co}})_{R-}$ and $(\dot{M}_{\mathrm{ac,tr,co}})_{R+}$, respectively, and the mass increasing rates from the surfaces $z=-z_{\mathrm{atm}}$ and $z=+z_{\mathrm{atm}}$ are added and denoted as $(\dot{M}_{\mathrm{ac,tr,co}})_{z}$. The signs of $(\dot{M}_{\mathrm{ac,tr,co}})_{R-}$, $(\dot{M}_{\mathrm{ac,tr,co}})_{R+}$, and $(\dot{M}_{\mathrm{ac,tr,co}})_{z}$ are determined so that they are positive when the total mass inside the enclosed region increases owing to its contribution. Since the surface at $R=R_{\mathrm{MRI}}$ ($R=R_{\eta}$) is shared between the active and transition zones (the transition and coherent zones), $(\dot{M}_{\mathrm{ac}})_{R+}=-(\dot{M}_{\mathrm{tr}})_{R-}$ ($(\dot{M}_{\mathrm{tr}})_{R+}=-(\dot{M}_{\mathrm{co}})_{R-}$) is satisfied. Figures 26a, 26c, and 26d show the time evolution of $(\dot{M}_{\mathrm{ac,tr,co}})_{R+}$, $(\dot{M}_{\mathrm{ac,tr,co}})_{R-}$ , $(\dot{M}_{\mathrm{ac,tr,co}})_{z}$, and $(\dot{M}_{\mathrm{ac,tr,co}})_{\mathrm{tot}}$.